A student scores 79 on an exam containing 100 problems. He tried to solve all the problems. His score was figured by subtracting two times the number of wrong answers from the number of correct answers. How many problems did he answer correctly?
Answer by checkley71(8403)

X-2(100-X)=79
X-200+2X=79
3X=79+200
3X=279
X=279/3
X=93 CORRECT ANSWERS.
100-93=7 WRONG ANSWERS.
93-2*7=93-14=79